OpenStudy (anonymous):

(x + 3) + 2x = 42 3x + 3 = 42 How did the first step get to the second one using the distribution property?

Mehek (mehek14):

multiplied by 1 \(\bf{1*(x+3)+2x=42\\1x+3+2x=42\\3x+3=42}\)

OpenStudy (anonymous):

I see! Thank you! What threw me off was the book answer didn't count "Combine Like Terms" as one of the steps. Thank you!

Mehek (mehek14):

yw ^_^